OpenStudy (anonymous):

A biologist is recording the loss of fish in a pond. She notes the number of fish, f, in the pond on June 1. On July 1 there were 63 fish in the pond, which is 52 fewer fish that were in the pond on June 1. Which equation represents this situation? A. f - 52 = 63 B. f - 63 = 52 C. 63 - f = 52 D. f - 63 = -52

OpenStudy (anonymous):

someone help please ;) i will give best response

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

The third sentence restated: 63 is 52 fewer than f. which means it is f-52. so, 63 = f - 52 That's the same as f - 52 = 63.

OpenStudy (anonymous):

^

OpenStudy (anonymous):

so whats my answer?

OpenStudy (anonymous):

A.